OpenStudy (anonymous):

Solve the following equation. x^2 + 7x = 0

OpenStudy (anonymous):

x=-7

jimthompson5910 (jim_thompson5910):

x^2 + 7x = 0 x(x+7) = 0 x=0 or x+7=0 x=0 or x=-7

OpenStudy (anonymous):

or x = 0

OpenStudy (anonymous):

hai guys!

jimthompson5910 (jim_thompson5910):

hello